Grade 11General Physics

Velocity of a train increases at a constant rate f1 from 0 to v and remains constant for an interval and finally decreases to 0at f2.
If x=total distance. Prove that the total time taken is x/v+v/2(1/f1+1/f2).

Nivashini Elango
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Let the train accelerate from the velocity 0 to v at this rate α in time t1 and distance is d1.
 f1 = v-0/t1 = v/t1
t1 = v/ f1
then d 1= ½ * ( )^2
                               = v^2/ 2 f1
After that it moves with constant speed v in time t2
Then d2 = vt2
Then it decelerates from v to zero in time t3 with the rate f2
  • t3 = v/ f2
No w using
  • v2+ u2 = 2 f2 d3
  • 0+ v2 = 2 f2 d3
  • d3 = v2/2 f2
Total distance (d) = v^2/ 2 α + vt2 + v2/2 f2
 d= v^2/ 2 f1 + vt2 + v2/2 f2
t2 = d/v- v/ 2 v/2
Then Total time
T = + d/v- v/ 2 v/2 + v/
    =d/v + v/2
As the total distance is 1
Then T = 1/v + v/2[1/f1 +1/f2].